### v1

@@ Abstract
The rapid integration of generative artificial intelligence (GenAI) technology into education necessitates precise measurement of GenAI literacy to ensure that learners and educators possess the skills to effectively engage with and critically evaluate this transformative technology. Existing instruments often rely on self-reports, which may be biased. In this study, we present the GenAI Literacy Assessment Test (GLAT), a 20-item multiple-choice instrument developed following established procedures in psychological and educational measurement. Structural validity and reliability were confirmed with responses from 355 higher education students using classical test theory and item response theory, resulting in a reliable 2-parameter logistic (2PL) model (Cronbach’s alpha = 0.80; omega total = 0.81) with a robust factor structure (RMSEA = 0.03; CFI = 0.97). Critically, GLAT scores were found to be significant predictors of learners’ performance in GenAI-supported tasks, outperforming self-reported measures such as perceived ChatGPT proficiency and demonstrating incremental validity. These results suggest that GLAT offers a reliable and more objective method for assessing GenAI literacy, with the potential to inform educational practices and policy decisions that aim to enhance learners’ and educators’ GenAI literacy, ultimately equipping them to navigate an AI-enhanced future.
@@ 1 Introduction
Generative artificial intelligence (GenAI) has rapidly emerged as a transformative force in higher education, challenging traditional pedagogical frameworks while simultaneously presenting novel opportunities for teaching, learning, and assessment. Tools like OpenAI’s ChatGPT, Google’s Gemini, and Anthropic’s Claude hold the potential to transform how personalised tutoring service can be delivered, how instructional materials can be generated, how lectures can be transcribed for accessibility, and how creativity can be nurtured through multimedia content generation (Bahroun et al., 2023; Yan et al., 2024a, b; Khosravi et al., 2023; Lee et al., 2023). However, the integration of these technologies is accompanied by complex challenges, including ethical considerations, risks of misinformation from model "hallucinations," and concerns regarding academic integrity (Ji et al., 2023; McDonald et al., 2024). Such complexities necessitate a deeper focus on fostering AI literacy, particularly GenAI literacy, among both educators and learners to fully harness GenAI’s benefits and mitigate its associated risks (Ng et al., 2021b; Zhao et al., 2024a).
@@
AI literacy refers to the set of competencies that enable individuals to effectively interact with AI technologies, encompassing understanding fundamental AI concepts, engaging in critical evaluation, and using AI tools ethically in diverse contexts (Long and Magerko, 2020; Ng et al., 2021b). Within this broad framework, GenAI literacy emerges as a specialised subset, focusing on skills required to engage with GenAI systems that can autonomously produce text, visuals, or other forms of media (Yan et al., 2024a; Annapureddy et al., 2024a). Developing GenAI literacy involves more than just foundational knowledge; it requires proficiency in crafting prompts, interpreting AI-generated outputs, and understanding the socio-ethical implications of using such tools (Zhao et al., 2024a; Bozkurt, 2024a). As GenAI becomes increasingly embedded in educational systems, it is imperative for learners and educators to acquire these competencies to effectively leverage the technology while minimising potential pitfalls such as biases or inaccuracies (Lyu et al., 2024b; Chiu, 2024a).
@@
Numerous instruments have been developed to assess AI literacy, reflecting the diversity of competencies that individuals need to navigate AI technologies. Conventional AI literacy assessments often rely on self-reported surveys, which are effective in capturing perceived knowledge but may lack the objectivity needed to accurately measure actual competencies, especially given the tendency for individuals to overestimate their understanding (Lintner, 2024; Laupichler et al., 2023b). Most existing instruments address general AI literacy, focusing on technical knowledge, awareness, and ethical considerations, but fail to adequately capture the unique skills required for GenAI (Koch et al., 2024; Zhao et al., 2024a). There is a growing demand for more nuanced and context-specific instruments to evaluate GenAI literacy, particularly as generative tools become integral to both physical and digital learning environments (Koch et al., 2024; Zhao et al., 2024a; Yan et al., 2024b).
@@
Current AI literacy assessments can be broadly categorised into two types: self-reported and performance-based measures. Self-reported instruments, while commonly used, provide insights into individuals’ perceived abilities but may introduce biases that obscure a more objective measure of literacy levels (Ng et al., 2021b; Lintner, 2024). In contrast, performance-based assessments evaluate actual competencies through direct engagement, offering a more reliable measure of skills. This distinction is especially pertinent for GenAI literacy, where there is often a gap between learners’ perceived understanding and their real ability to effectively utilise generative tools (Lyu et al., 2024b). GenAI technologies necessitate iterative, context-specific interactions that require both sophisticated prompting skills and the ability to critically assess AI outputs, areas where self-reports may fall short (Chiu, 2024a; Bozkurt, 2024a). Therefore, developing performance-based instruments is essential to provide an accurate assessment of individuals’ abilities to engage with these advanced technologies in educational settings. However, to the best of our knowledge, there are still limited performance-based tools for measuring students’ GenAI literacy in higher education, particularly those that have been rigorously developed and validated according to established psychological and educational measurement standards (Thorndike et al., 1991; American Educational Research Association et al., 2014).
@@
The current study contributes to the field of AI in education and AI literacy by introducing the GenAI Literacy Assessment Test (GLAT), a performance-based instrument specifically designed to evaluate GenAI literacy within higher education contexts. The GLAT aims to fill a critical gap in existing assessment tools by providing a more objective, comprehensive evaluation of the key competencies required to interact with GenAI tools. Unlike existing assessments that focus predominantly on general AI skills, GLAT targets the unique skills necessary for effective engagement with generative technologies, including technical proficiency, ethical awareness, and the capacity for critical evaluation of GenAI-generated outputs. This instrument is grounded in rigorous methodologies from psychological and educational measurement (Thorndike et al., 1991; American Educational Research Association et al., 2014), ensuring both validity and reliability. By focusing on performance-based metrics, GLAT provides educators and researchers with an accurate tool to assess how well students and educators understand and can leverage GenAI technologies, ultimately informing targeted interventions that can enhance these competencies.
@@ 2.1 AI and GenAI Literacy
The rapid advancements in AI technologies have accentuated the importance of AI literacy, especially in educational research. Researchers have endeavoured to define and conceptualise "AI literacy," with the definition by Long and Magerko (2020) being frequently cited: "a set of competencies that enables individuals to critically evaluate AI technologies; communicate and collaborate effectively with AI; and use AI as a tool online, at home, and in the workplace." Building upon this definition, other researchers have explored essential aspects of AI literacy. For instance, Kandlhofer et al. (2016) and Burgsteiner et al. (2016) concentrated on the comprehension of fundamental AI concepts present in various products and services. Meanwhile, Wang et al. (2023) emphasised critical evaluation, practical application, and ethical responsibilities. Additionally, Ng et al. (2021b, a) and Almatrafi et al. (2024) have refined this framework by stressing competencies such as recognition, application, evaluation, creation, and ethical navigation.
@@
The rise of GenAI technologies, like ChatGPT, necessitates a re-evaluation of AI literacy within the specific context of generative technologies. GenAI’s capability to generate substantial content from minimal input alters the landscape, prompting a need for a revised understanding of AI literacy in this context (Zhao et al., 2024b). Although there has been increasing academic interest in GenAI, a comprehensive definition of GenAI literacy remains elusive (Annapureddy et al., 2024b). Many current AI literacy frameworks are too general and do not address the specific competencies required by GenAI, which differ significantly from those of predictive models. Specifically, GenAI literacy calls for an integrative approach that combines theoretical knowledge, practical skills, and critical reflection. The 3wAI Framework by Bozkurt (2024b) addresses this need, focusing on "Know What," "Know How," and "Know Why"—aiming to promote foundational understanding, practical application, and ethical awareness. Zhao et al. (2024b) asserts that GenAI literacy should include pragmatic, safety, reflective, socio-ethical, and contextual elements. Scholars such as Lyu et al. (2024a) and Annapureddy et al. (2024b) emphasise the distinction between general AI literacy and the specific skills required for GenAI, pointing out that existing frameworks often overlook the skills necessary for effectively utilising these tools.
@@
The need for GenAI literacy is particularly crucial in educational settings. The absence of a comprehensive GenAI literacy framework poses challenges to its effective integration into learning environments (Annapureddy et al., 2024b). Chiu (2024b) underlines the necessity of empirically evaluating pedagogies that incorporate GenAI to determine their impact on student outcomes. GenAI literacy can significantly enhance language learning, as suggested by Alzubi (2024), while Bozkurt (2023) advocates for its inclusion in curricula to prepare students for an AI-augmented future. Despite the potential of GenAI tools to improve student learning, Lyu et al. (2024a) discovered that student-generated prompts often lack quality, highlighting a deficiency in necessary skills. This underscores the urgent need to develop GenAI literacy so that students can fully harness the potential of these technologies.
@@ 2.2 AI Literacy Instruments
Multiple AI literacy measurement instruments have been developed to address various contexts, audiences, and facets of AI literacy. These facets encompass technical, ethical, behavioural, and contextual elements, reflecting the multifaceted nature of AI literacy. The Scale for the Assessment of Non-Experts’ AI Literacy (SNAIL), created by Laupichler et al. (2023a), evaluates technical knowledge, critical analysis, and practical application of AI. This scale’s validity was confirmed through factor analyses and a Delphi study (Laupichler et al., 2023b). The AI Literacy Scale (AILS) focuses on general AI literacy by measuring awareness, usage, evaluation, and ethics and was validated by subject matter experts (Wang et al., 2023). For specific audiences, the Medical Artificial Intelligence Readiness Scale for Medical Students (MAIRS-MS) targets medical students, assessing cognition, ability, vision, and ethics (Karaca et al., 2021). Another survey, based on the Unified Theory of Acceptance and Use of Technology (UTAUT) and the Technological Pedagogical and Content Knowledge (TPACK) framework, examines pedagogical knowledge and AI use intentions among EFL teachers (An et al., 2023). Instruments for younger audiences include the AI Literacy Questionnaire (AILQ) by Ng et al.2024, designed for secondary students, which assesses affective, behavioural, cognitive, and ethical dimensions. Chai et al.’s AI Literacy Instrument 2021 explores students’ confidence, readiness, and perceptions of AI. For broader competencies, Carolus et al.’s Meta AI Literacy Scale (MAILS) 2023 covers ethics, persuasion literacy, and emotion regulation. Additionally, Pinski et al.’s AI Literacy Instrument 2023 targets AI professionals, focusing on human-AI interaction, AI processing, and task knowledge, with validation for reliability and robustness. Lastly, Lee and Park’s ChatGPT Literacy Tool 2024 is currently the only instrument specifically designed to assess GenAI skills among university students, though it relies on self-reported data.
@@
Despite progress in AI literacy assessment, substantial gaps persist. Most current instruments rely heavily on self-reported assessments, with minimal use of performance-based measurements. A systematic literature review by Lintner (2024) identified 13 self-reported and only three performance-based instruments, highlighting the predominant reliance on self-reported tools and the urgent need for more objective assessments. While there are some performance-based measures for general AI literacy, such as Hornsberger’s 2023 test, which includes 30 multiple-choice questions and a sorting item, and Chiu’s 2024 test consisting of 25 multiple-choice questions, no such measures have been developed for GenAI literacy. Performance-based evaluations are crucial for GenAI literacy, given the importance of practical engagement and iterative interactions (Lintner, 2024; Laupichler et al., 2023b; Yan et al., 2024a). To address this gap, it is essential to develop new GenAI literacy instruments specifically targeting GenAI skills and incorporating performance-based assessments for a more accurate evaluation of learners’ competencies. These instruments would help educators better understand how students interact with generative models, identify areas needing additional training, and design effective interventions to enhance students’ abilities to use GenAI technologies.
@@ 2.4 Generative AI Literacy Assessment Test (GLAT)
Building on the foundations of CTT and IRT, this study introduces the development and validation of the GLAT. Designed to measure GenAI literacy among higher education students, the GLAT stands on the robust foundations of psychological and educational measurement practices (Thorndike et al., 1991; American Educational Research Association et al., 2014). These foundations ensure that the GLAT can effectively assess key aspects of GenAI literacy, including foundational knowledge, application, ethical awareness, and critical evaluation capabilities (Annapureddy et al., 2024a; Long and Magerko, 2020). Specifically, the following research questions were investigated, aiming to assess the validity and reliability of the GLAT in capturing learners’ GenAI literacy and predicting their learning performance:
@@
  • RQ2: To what extent does the GLAT measure of GenAI literacy predict learners’ performance in tasks with GenAI chatbots compared to self-reported instruments?

### v2

@@ Abstract
The rapid integration of generative artificial intelligence (GenAI) technology into education necessitates precise measurement of GenAI literacy to ensure that learners and educators possess the skills to engage with and critically evaluate this transformative technology effectively. Existing instruments often rely on self-reports, which may be biased. In this study, we present the GenAI Literacy Assessment Test (GLAT), a 20-item multiple-choice instrument developed following established procedures in psychological and educational measurement. Structural validity and reliability were confirmed with responses from 355 higher education students using classical test theory and item response theory, resulting in a reliable 2-parameter logistic (2PL) model (Cronbach’s alpha = 0.80; omega total = 0.81) with a robust factor structure (RMSEA = 0.03; CFI = 0.97). Critically, GLAT scores were found to be significant predictors of learners’ performance in GenAI-supported tasks, outperforming self-reported measures such as perceived ChatGPT proficiency and demonstrating external validity. These results suggest that GLAT offers a reliable and valid method for assessing GenAI literacy, with the potential to inform educational practices and policy decisions that aim to enhance learners’ and educators’ GenAI literacy, ultimately equipping them to navigate an AI-enhanced future.
@@ 1 Introduction
Generative artificial intelligence (GenAI) has rapidly emerged as a transformative force in higher education, challenging traditional pedagogical frameworks while simultaneously presenting novel opportunities for teaching, learning, and assessment. Tools like OpenAI’s ChatGPT, Google’s Gemini, and Anthropic’s Claude hold the potential to transform how personalised tutoring service can be delivered, how instructional materials can be generated, how lectures can be transcribed for accessibility, and how creativity can be nurtured through multimedia content generation (Yan et al., 2024a; Khosravi et al., 2023). However, the integration of these technologies is accompanied by complex challenges, including ethical considerations, risks of misinformation from model "hallucinations," and concerns regarding academic integrity (Ji et al., 2023; McDonald et al., 2024). Such complexities necessitate a deeper focus on fostering AI literacy, particularly GenAI literacy, among both educators and learners to fully harness GenAI’s benefits and mitigate its associated risks (Ng et al., 2021b; Zhao et al., 2024).
@@
AI literacy refers to the set of competencies that enable individuals to effectively interact with AI technologies, encompassing understanding fundamental AI concepts, engaging in critical evaluation, and using AI tools ethically in diverse contexts (Long and Magerko, 2020; Ng et al., 2021b). Within this broad framework, GenAI literacy emerges as a specialised subset, focusing on skills required to engage with GenAI systems that can autonomously produce text, visuals, or other forms of media (Yan et al., 2024a; Annapureddy et al., 2024). Developing GenAI literacy involves more than just foundational knowledge; it requires proficiency in crafting prompts, interpreting AI-generated outputs, and understanding the socio-ethical implications of using such tools (Zhao et al., 2024; Bozkurt, 2024a). As GenAI becomes increasingly embedded in educational systems, it is imperative for learners and educators to acquire these competencies to effectively leverage the technology while minimising potential pitfalls such as biases or inaccuracies (Lyu et al., 2024; Chiu, 2024).
@@
Numerous instruments have been developed to assess AI literacy, reflecting the diversity of competencies that individuals need to navigate AI technologies. Conventional AI literacy assessments often rely on self-reported surveys, which are effective in capturing perceived knowledge but may lack the reliability needed to accurately measure actual competencies, especially given the tendency for individuals to overestimate their understanding (Lintner, 2024; Laupichler et al., 2023b). Most existing instruments address general AI literacy, focusing on technical knowledge, awareness, and ethical considerations, but fail to adequately capture the unique skills required for GenAI (Koch et al., 2024; Zhao et al., 2024). There is a growing demand for more nuanced and context-specific instruments to evaluate GenAI literacy, particularly as generative tools become integral to both physical and digital learning environments (Koch et al., 2024; Zhao et al., 2024; Yan et al., 2024b).
@@
Current AI literacy assessments can be broadly categorised into two types: self-reported and performance-based measures. Self-reported instruments, while commonly used, provide insights into individuals’ perceived abilities but may introduce biases that obscure a more reliable measure of literacy levels (Ng et al., 2021b; Lintner, 2024). In contrast, performance-based assessments evaluate actual competencies through direct engagement, offering a more reliable measure of skills. This distinction is especially pertinent for GenAI literacy, where there is often a gap between learners’ perceived understanding and their real ability to effectively utilise generative tools (Lyu et al., 2024). GenAI technologies necessitate iterative, context-specific interactions that require both sophisticated prompting skills and the ability to critically assess AI outputs, areas where self-reports may fall short (Chiu, 2024; Bozkurt, 2024a). Therefore, developing performance-based instruments is essential to provide a reliable assessment of individuals’ abilities to engage with these advanced technologies in educational settings. However, to the best of our knowledge, there are still limited performance-based tools for measuring students’ GenAI literacy in higher education, particularly those that have been rigorously developed and validated according to established psychological and educational measurement standards (Thorndike et al., 1991; American Educational Research Association et al., 2014).
@@
The current study contributes to the field of AI in education and AI literacy by introducing the GenAI Literacy Assessment Test (GLAT), a performance-based instrument specifically designed to evaluate GenAI literacy within higher education contexts. The GLAT aims to fill a critical gap in existing assessment tools by providing a more reliable, comprehensive evaluation of the key competencies required to interact with GenAI tools. Unlike existing assessments that focus predominantly on general AI skills, GLAT targets the unique skills necessary for effective engagement with generative technologies, including technical proficiency, ethical awareness, and the capacity for critical evaluation of GenAI-generated outputs. This instrument is grounded in rigorous methodologies from psychological and educational measurement (Thorndike et al., 1991; American Educational Research Association et al., 2014), ensuring both validity and reliability. By focusing on performance-based metrics, GLAT provides educators and researchers with a reliable tool to assess how well students and educators understand and can leverage GenAI technologies, ultimately informing targeted interventions that can enhance these competencies.
@@ 2.1 AI and GenAI Literacy
The rapid advancements in AI technologies have accentuated the importance of AI literacy, especially in educational research. Researchers have endeavoured to define and conceptualise "AI literacy," with the definition by Long and Magerko (2020, p.2) being frequently cited: "a set of competencies that enables individuals to critically evaluate AI technologies; communicate and collaborate effectively with AI; and use AI as a tool online, at home, and in the workplace." Building upon this definition, other researchers have explored essential aspects of AI literacy. For instance, Kandlhofer et al. (2016) and Burgsteiner et al. (2016) concentrated on the comprehension of fundamental AI concepts present in various products and services. Meanwhile, Wang et al. (2023) emphasised critical evaluation, practical application, and ethical responsibilities. Additionally, Ng et al. (2021b), Ng et al. (2021a), and Almatrafi et al. (2024) have refined this framework by stressing competencies such as recognition, application, evaluation, creation, and ethical navigation.
@@
The rise of GenAI technologies, like ChatGPT, necessitates a re-evaluation of AI literacy within the specific context of generative technologies. GenAI’s capability to generate substantial content from minimal input alters the landscape, prompting a need for a revised understanding of AI literacy in this context (Zhao et al., 2024). Although there has been increasing academic interest in GenAI, a comprehensive definition of GenAI literacy remains elusive (Annapureddy et al., 2024). Many current AI literacy frameworks are too general and do not address the specific competencies required by GenAI, which differ significantly from those of predictive models. Specifically, GenAI literacy calls for an integrative approach that combines theoretical knowledge, practical skills, and critical reflection. The 3wAI Framework by Bozkurt (2024b) addresses this need, focusing on "Know What," "Know How," and "Know Why" and aiming to promote foundational understanding, practical application, and ethical awareness. Zhao et al. (2024) asserts that GenAI literacy should include pragmatic, safety, reflective, socio-ethical, and contextual elements. Scholars such as Lyu et al. (2024) and Annapureddy et al. (2024) emphasise the distinction between general AI literacy and the specific skills required for GenAI, pointing out that existing frameworks often overlook the skills necessary for effectively utilising these tools.
@@
The need for GenAI literacy is particularly crucial in educational settings. The absence of a comprehensive GenAI literacy framework poses challenges to its effective integration into learning environments (Annapureddy et al., 2024). Chiu (2024) underlines the necessity of empirically evaluating pedagogies that incorporate GenAI to determine their impact on student outcomes. GenAI literacy can significantly enhance language learning, as suggested by Alzubi (2024), while Bozkurt (2023) advocates for its inclusion in curricula to prepare students for an AI-augmented future. Despite the potential of GenAI tools to improve student learning, Lyu et al. (2024) discovered that student-generated prompts often lack quality, highlighting a deficiency in necessary skills. This underscores the urgent need to develop GenAI literacy so that students can fully harness the potential of these technologies.
@@ 2.2 AI Literacy Instruments
Multiple AI literacy measurement instruments have been developed to address various contexts, audiences, and facets of AI literacy. These facets encompass technical, ethical, behavioural, and contextual elements, reflecting the multifaceted nature of AI literacy. The Scale for the Assessment of Non-Experts’ AI Literacy (SNAIL), created by Laupichler et al. (2023a), evaluates technical knowledge, critical analysis, and practical application of AI. This scale’s validity was confirmed through factor analyses and a Delphi study (Laupichler et al., 2023b). The AI Literacy Scale (AILS) focuses on general AI literacy by measuring awareness, usage, evaluation, and ethics and was validated by subject matter experts (Wang et al., 2023). For specific audiences, the Medical Artificial Intelligence Readiness Scale for Medical Students (MAIRS-MS) targets medical students, assessing cognition, ability, vision, and ethics (Karaca et al., 2021). Another survey, based on the Unified Theory of Acceptance and Use of Technology (UTAUT) and the Technological Pedagogical and Content Knowledge (TPACK) framework, examines pedagogical knowledge and AI use intentions among EFL teachers (An et al., 2023). Instruments for younger audiences include the AI Literacy Questionnaire (AILQ) by Ng et al.2024, designed for secondary students, which assesses affective, behavioural, cognitive, and ethical dimensions. Chai et al.’s 2021 AI Literacy Instrument explores students’ confidence, readiness, and perceptions of AI. For broader competencies, Carolus et al.’s 2023 Meta AI Literacy Scale (MAILS) covers ethics, persuasion literacy, and emotion regulation. Additionally, Pinski et al.’s 2023 AI Literacy Instrument targets AI professionals, focusing on human-AI interaction, AI processing, and task knowledge, with validation for reliability and robustness. Lastly, Lee and Park’s 2024 ChatGPT Literacy Tool is currently the only instrument specifically designed to assess GenAI skills among university students, though it relies on self-reported data.
@@
Despite progress in AI literacy assessment, substantial gaps persist. Most current instruments rely heavily on self-reported assessments, with minimal use of performance-based measurements. A systematic literature review by Lintner (2024) identified 13 self-reported and only three performance-based instruments, highlighting the predominant reliance on self-reported tools and the urgent need for more reliable assessments. While there are some performance-based measures for general AI literacy, such as Hornsberger’s 2023 test, which includes 30 multiple-choice questions and a sorting item, and Chiu’s 2024 test consisting of 25 multiple-choice questions, no such measures have been developed for GenAI literacy. Performance-based evaluations are crucial for GenAI literacy, given the importance of practical engagement and iterative interactions (Lintner, 2024; Laupichler et al., 2023b; Yan et al., 2024a). To address this gap, it is essential to develop new GenAI literacy instruments specifically targeting GenAI skills and incorporating performance-based assessments for a more accurate evaluation of learners’ competencies. These instruments would help educators better understand how students interact with generative models, identify areas needing additional training, and design effective interventions to enhance students’ abilities to use GenAI technologies.
@@ 2.3 Classical Test Theory and Item Response Theory
2.4 External Validity and Domain Knowledge
External validity is a crucial aspect of educational and psychological assessments, ensuring that the constructs measured by an instrument, such as GenAI literacy, can predict relevant learning outcomes and performances in varied contexts (Messick, 1995; Boateng et al., 2018). This concept is emphasised in Messick’s unified theory of validity, which posits that the validation process must consider not only how well an instrument measures the intended construct but also how well the construct aligns with real-world tasks and external criteria (Messick, 1995). In the context of GenAI literacy, external validity involves the instrument’s ability to accurately predict students’ capacity to engage with and perform learning tasks using GenAI tools effectively. This predictive capability is critical, as the ultimate goal of assessing GenAI literacy is to ensure that learners are equipped with the skills necessary to apply these technologies in authentic educational and professional settings (Yan et al., 2024a; Annapureddy et al., 2024).
Domain knowledge plays a vital role in assessing the external validity of GenAI literacy tools (Alexander, 1992; Alexander and Judy, 1988). It is essential to consider and control for the varying levels of domain knowledge possessed by students (Alexander and Judy, 1988), as this can significantly influence how effectively they can utilise GenAI technologies within different learning tasks (Tricot and Sweller, 2014). Failure to account for domain knowledge could result in skewed assessments (Alexander, 1992), where the lack of subject-specific understanding might be mistaken for deficiencies in GenAI literacy itself. Consequently, assessing GenAI literacy necessitates a nuanced approach that considers the interplay between domain knowledge and the specific skills required to interact with GenAI tools.
@@ 2.5 Generative AI Literacy Assessment Test (GLAT)
Building on these theoretical foundations, this study introduces the development and validation of the GLAT. Designed to measure GenAI literacy among higher education students, the GLAT stands on the robust foundations of psychological and educational measurement practices (Thorndike et al., 1991; American Educational Research Association et al., 2014). These foundations ensure that the GLAT can effectively assess key aspects of GenAI literacy, including foundational knowledge, application, ethical awareness, and critical evaluation capabilities (Annapureddy et al., 2024; Long and Magerko, 2020). Specifically, the following research questions were investigated, aiming to assess the validity and reliability of the GLAT in capturing learners’ GenAI literacy and predicting their learning performance:
@@
  • RQ2: To what extent does the GLAT measure of GenAI literacy demonstrate external validity by predicting learners’ performance in learning tasks with GenAI chatbots compared to self-reported instruments?